OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (anonymous):

You know the slope/gradient of the line has to be the same (because the 2 lines are parallel), so m =-2/3. You then simply insert the coordinates of the point and m into the formula y-y1=m(x-x1), where y1 and x1 are the coordinates. You therefore get y+3=-2/3(x-9) now you just rearrange the formula to get y=(-2/3)x+3